# From Prejudice to Parity: A New Approach to Debiasing Large Language Model Word Embeddings

### v2

@@ 5.2 Stereotype Score
Our research focuses on evaluating and mitigating stereotypical bias in natural language inference tasks using the Stereoset dataset. This dataset comprises pairs of sentences differing only in the substitution of words related to social groups like gender, race, or religion. The objective is to predict their relationship as same, entailment, or contradiction. We introduce a method aimed at reducing bias in word embeddings, with $SS$ values closer to 50 indicating decreased bias. Table 2 presents DeepSoftDebias’s results alongside existing approaches on the Stereoset dataset. Notably, DeepSoftDebias achieves the lowest $SS$ across all social groups, demonstrating its effectiveness in bias reduction. Particularly impressive is DeepSoftDebias’s performance in the gender and race categories, where it significantly outperforms existing methods. For instance, with the SFR-Embedding-Mistral Jiang et al. (2023) model, DeepSoftDebias achieves an $SS$ of 50 for gender and 50.409 for race using the Llama-2-7b model. Additionally, DeepSoftDebias attains a score of 51.282 (or 48.717) for the Zephyr-7b-beta Tunstall et al. (2023) (or Alpaca-7b Taori et al. (2023)) group. We present an illustration of these scores in Fig. 6.
@@ 5.3 Crows-Pairs Dataset
where stereoscore is the number of stereotypical samples that agree with their label direction and antistereoscore is the number of anti-stereotypical samples that agree with their label direction. Particularly noteworthy is DeepSoftDebias’s superior performance in the gender and religion categories. For instance, with the Yi-6B model, DeepSoftDebias achieves a score of 49.62 for gender and 50.48 for religion with the falcon-7b model. Similarly, using the SFR-Embedding-Mistral model, DeepSoftDebias achieves a score of 49.03 for race biasing the SFR-Embedding-Mistral model. These results underscore the effectiveness of DeepSoftDebias in mitigating social bias in word embeddings. We depict the variation of these scores in Fig. 8.
@@
6 DownStream Results
@@
6.1 Sentiment Classification
In our study, we employ downstream testing to assess the utility of embeddings debiased using DeepSoftDebias across two key natural language processing tasks: text classification and named entity recognition (NER). Utilizing the IMDB Sentiment Classification dataset İlhan Tarımer et al. (2019) for text classification, featuring labeled movie reviews as positive or negative, we compute the average sentence vectors using both original and debiased embeddings. Training XGBoost Chen and Guestrin (2016) classifiers on these vectors, we compare their accuracy on the test set, recognizing accuracy as a straightforward metric for binary classification tasks like sentiment analysis. Notably, our results reveal a performance improvement when debiasing in the gender and religion directions, whereas a slight decrease in performance is observed in the case of race debiasing. We provide a visual representation of these results in Fig. 9.
6.2 Named Entity Recognition (NER)
In our research, we examine the performance of debiased embeddings in the domain of named entity recognition (NER) using the Reuters subset of the CrossNER Liu et al. (2020) dataset. This dataset comprises news domain sentences annotated with four entity types: person, location, organization, and product. Employing a simple BiLSTM model, we input padded arrays of embeddings for each sentence and trained the model on the dataset. We evaluate the models’ performance on the test set using the macro-averaged F1-score, a metric that balances precision and recall, crucial for accurate entity identification and classification. To mitigate potential bias towards more frequent entity types, we adopt macro-averaging, allotting equal importance to each entity type. Remarkably, our findings indicate a slight performance boost when using debiased embeddings in all three directions compared to biased embeddings. We provide a visual representation of these results in Fig. 10.
7 Ablation Experiments
In our study, we conduct ablation experiments to assess the effectiveness of various debiasing techniques in the realm of natural language processing. These techniques encompassed five distinct scenarios: the utilization of debiased embeddings, the application of the original soft debiasing method, the integration of the original debiasing method with the Adam optimizer, the incorporation of new neural networks trained with the debiasing method alongside the SGD optimizer, and finally, our proposed approach, which combines neural networks trained with the debiasing method and the Adam optimizer.
Through rigorous experimentation across three biasing directions, we systematically analyze the performance of each method. Our results reveal a consistent trend of incremental improvements as we transitioned from one method to the next. Notably, DeepSoftDebias, emerged as the standout performer, boasting the highest mean average cosine similarity score across all evaluated scenarios. In addition, our analysis revealed that substituting the transformation matrix with our neural network approach resulted in the most significant enhancement in the efficacy of the debiasing method. This observation underscores the pivotal role played by neural networks in maximizing the effectiveness of the debiasing techniques. We present a visualization of the results of our Ablation experiments in Table 5.
8 Discussion
In this section, we summarise the answers to our research questions.
8.1 RQ1
8.2 RQ2
As a step for future work, we are curious to investigate scaling patterns to a further extent.
8.3 RQ3
8.4 RQ4
9 Conclusion
In this paper, we propose DeepSoftDebias, an approach that leverages neural networks to reduce bias in large language model embeddings. We perform an exhaustive series of tests using multiple performance metrics, state-of-the-art datasets, and downstream tasks to ensure that our debiasing technique is robust, efficient, and accurate. In the future, it would be interesting to see how this method translates to multilingual datasets since bias is language and culture-specific. We hope that this research paves the way for future endeavors that look to make large language models fair, ethical, and bias-free.
10 Limitations
While we do perform exhaustive analysis to test our proposed methodology, our study is monolingual and covers datasets only in English. Consequently, our downstream tasks are also tested only in English. Further, we were unable to test on API-based models at this time.
11 Ethics Statement
Appendix
Appendix A MAC Scores of DeepSoftDebias
Appendix B Stereoset Scores of DeepSoftDebias
Appendix C Downstream Testing Results

### v4

@@ 6.2 Stereotype Score
Our research focuses on evaluating and mitigating stereotypical bias in NLI tasks using the Stereoset dataset. This dataset comprises pairs of sentences differing only in the substitution of words related to social groups like gender, race, or religion. The objective is to predict their relationship as same, entailment, or contradiction. We introduce a method aimed at reducing bias in word embeddings, with Stereotype Score $SS$ values closer to 50 indicating decreased bias. Table 2 presents DeepSoftDebias’s results alongside existing approaches on the Stereoset dataset. Notably, DeepSoftDebias achieves the lowest $SS$ across all social groups, demonstrating its effectiveness in bias reduction. Particularly impressive is DeepSoftDebias’s performance in the gender and race categories, where it significantly outperforms existing methods. For instance, with the SFR-Embedding-Mistral Jiang et al. (2023) model, DeepSoftDebias achieves an $SS$ of 50 for gender and 50.409 for race using the Llama-2-7b model. Additionally, DeepSoftDebias attains a score of 51.282 for the Zephyr-7b-beta Tunstall et al. (2023) or 48.717 for Alpaca-7b Taori et al. (2023). We present these score in 2 an illustration of these scores in Fig. 6.
@@ 6.3 Crows-Pairs Dataset
where Crows Pair Stereotype Score (CSS) is the number of stereotypical samples that agree with their label direction and Crows Pairs Anti-stereotype Score (CAS) is the number of anti-stereotypical samples that agree with their label direction. Label direction refers to the label given the pair of sentences whether they are stereotypical or anti-stereotypical. In our evaluation we get the average sentence vector of the context and the more and less (anti-)stereotypical sentence. We then see whether the context vector is closer to the more (anti-)stereotypical sentence or the less (anti-)stereotypical sentence. If it is closer to the more (anti–)stereotypical sentence,, then we state that it agrees with the (anti–)stereotype, i.e., the label direction. Particularly noteworthy is DeepSoftDebias’s superior performance in the gender and religion categories. For instance, with the Yi-6B model, DeepSoftDebias achieves a score of 49.62 for gender and 50.48 for religion with the falcon-7b model. Similarly, using the SFR-Embedding-Mistral model, DeepSoftDebias achieves a score of 49.03 for race biasing the SFR-Embedding-Mistral model. These results underscore the effectiveness of DeepSoftDebias in mitigating social bias in word embeddings. We present these score in 3 and depict the variation of these scores in Fig. 8.
We also report the CSS and CAS score which refer to the CrossNER Stereotype score, i.e., the number of times the model agrees with the more stereotypes statement when the label direction is stereotype, and the CrossNER Anti-stereotype score, which refers to the number of times the model agrees with the more anti-stereotyped statement when the label direction was anti-sterotype.
@@
7 Discussion
In this section, we summarise the answers to our research questions.
RQ1
RQ2
As a step for future work, we are curious to investigate scaling patterns to a further extent.A visualization of this is provided in Fig 11
RQ3
RQ4
8 Conclusion
In this paper, we propose DeepSoftDebias, an approach that leverages neural networks to reduce bias in large language model embeddings. We perform an exhaustive series of tests using multiple performance metrics, state-of-the-art datasets, and downstream tasks to ensure that our debiasing technique is robust, efficient, and accurate. In the future, it would be interesting to see how this method translates to multilingual datasets since bias is language and culture-specific. We hope that this research paves the way for future endeavors that look to make LLMs fair, ethical, and bias-free.
9 Limitations
While we do perform exhaustive analysis to test our proposed methodology, our study is monolingual and covers datasets only in English. Consequently, our downstream tasks are also tested only in English. Further, we were unable to conduct test on API-based models at this time. Our testing was also constrained by the limitations of GPU VRAM, which prevented us from extending our testing to larger models such as Llama-65B. These models could not be accommodated within the GPU VRAM, even after applying quantization to 8 bits. Consequently, the largest model that we were able to test was the Gemma-7B model.
10 Ethics Statement
Appendix
Appendix A Table of words and bias they represent
Bias Direction Biased Words
Gender Male "manager", "executive", "doctor", "lawyer", "programmer", "scientist", "soldier", "supervisor", "rancher", "janitor", "firefighter", "officer"
Female "secretary", "nurse", "clerk", "artist", "homemaker", "dancer", "singer", "librarian", "maid", "hairdresser", "stylist", "receptionist", "counselor"
Race Black "slave", "musician", "runner", "criminal", "homeless"
Asian "manager", "executive", "redneck", "hillbilly", "leader", "farmer"
Caucasian "doctor", "engineer", "laborer", "teacher"
Religion Jew "greedy", "cheap", "hairy", "liberal"
Christian "judgemental", "conservative", "familial"
Muslim "violent", "terrorist", "dirty", "uneducated"
Table 4: List of Words related to sub-categories in the bias directions explored
Appendix B MAC Scores of DeepSoftDebias
Appendix C Stereoset Scores of DeepSoftDebias
Appendix D Downstream Testing Results
Model Variant Baseline Debiased Text Class. Acc. DSB Debiased Text Class. Acc. Baseline Debiased NER Macro F1 Avg. DSB Debiased NER Macro F1 Avg.
Gemma gemma-2b 0.7655 0.7964 0.469 0.484
BAAI bge-base-en-v1.5 0.8296 0.822 0.458 0.421
Mistral SFR-Embedding-Mistral 0.8297 0.821 0.404 0.428
Gemma gemma-7b 0.516 0.8032 0.198 0.475
Zephyr zephyr-7b-beta 0.7997 0.81 0.403 0.429
mxbai mxbai-embed-large-v1 0.8366 0.7903 0.461 0.455
Table 5: Downstream testing results comparison with embeddings debiased using DeepSoftDebias and the baseline SoftDebais Method. The first two columns represent results for downstream performance on sentiment analysis. The second two columns represent results for downstream performance on NER.
In our research, we primarily focus on the debiasing of word embeddings derived from Language Learning Models (LLMs). We aim to investigate the impact of this debiasing on the performance of these embeddings when subjected to identical training and testing methodologies. Our objective is to quantitatively measure any performance fluctuations (increase or decrease) on the downstream tasks that we test. For this purpose, we trained simple models on top of these word embeddings. For instance, we used an XGBoost model without any hyperparameter tuning for the classification task, and a straightforward bidirectional LSTM for the Named Entity Recognition (NER) task. It is important to note that our goal in presenting our results on these two tasks is not to establish a benchmark for debiased embeddings. Instead, we aim to demonstrate the effect of debiasing on the performance of word embeddings in downstream tasks, as seen in the seminal work of Gonen and Goldberg (2019). This approach allows us to provide a more comprehensive understanding of the implications and potential benefits of debiasing word embeddings.
D.1 Sentiment Classification
In our study, we employ downstream testing to assess the utility of embeddings debiased using DeepSoftDebias across two key natural language processing tasks: text classification and named entity recognition (NER). Utilizing the IMDB Sentiment Classification dataset İlhan Tarımer et al. (2019) and Stanford TreeBank Dataset for text classification, featuring labeled movie reviews as positive or negative, we compute the average sentence vectors using both original and debiased embeddings. Training XGBoost Chen and Guestrin (2016) classifiers on these vectors, we compare their accuracy on the test set, recognizing accuracy as a straightforward metric for binary classification tasks like sentiment analysis. Notably, our results reveal a performance improvement when debiasing in the gender and religion directions, whereas a slight decrease in performance is observed in the case of race debiasing. We provide these results in Table:6 for IMDB Sentiment classification and Table:7 for Stanford Sentiment Treebank. A visual representation of these results in Fig. 9.
D.2 Named Entity Recognition (NER)
In our research, we examine the performance of debiased embeddings in the domain of named entity recognition (NER) using the Reuters subset of the CrossNER Liu et al. (2020) dataset. This dataset comprises news domain sentences annotated with four entity types: person, location, organization, and product. Employing a simple BiLSTM model, we input padded arrays of embeddings for each sentence and trained the model on the dataset. We evaluate the models’ performance on the test set using the macro-averaged F1-score, a metric that balances precision and recall, crucial for accurate entity identification and classification. To mitigate potential bias towards more frequent entity types, we adopt macro-averaging, allotting equal importance to each entity type. Remarkably, our findings indicate a slight performance boost when using debiased embeddings in all three directions compared to biased embeddings. We provide these results in Table:6and a visual representation of these results in Fig. 10.
D.3 Semantic Textual Similarity
In our research, we evaluate the performance of debiased embeddings for the Semantic Textual Similarity (STS) task using the STS-B dataset. This dataset, a component of the General Language Understanding Evaluation (GLUE) benchmark, is a valuable resource for the STS task. The task aims to quantify the semantic similarity between two sentences, assigning a score from 1 to 5 based on their degree of semantic equivalence. The STS-B dataset, comprising examples from diverse sources, includes human annotations for sentence pair similarity, contributing significantly to the broader field of natural language understanding by facilitating the measurement of meaning equivalence across sentences. To utilize the embeddings for the task, we train a dual-head neural network. We perform cosine similarity after passing the average sentence vector of the two sentences through the network, followed by a Fully Connected layer to obtain the actual score. The performance of our approach is evaluated using Pearson’s correlation and Spearman’s correlation as metrics. This methodology allows us to develop and evaluate models’ ability to understand nuanced semantic relationships in text effectively. We provide our results in this task in Table:7
@@ D.3 Semantic Textual Similarity
Model Name Topic STS-B $\uparrow$ Baseline Debiased PCC STS-B $\uparrow$ DeepSoftDebias Debiased PCC SST Biased Acc. SST Baseline Debiased Acc. SST DeepSoftDebias Debiased Acc.
BAAI/bge-base-en-v1.5 Gender 0.088 0.001 0.730 0.725 0.693
BAAI/bge-large-en-v1.5 0.159 0.105 0.727 0.710 0.705
google/gemma-2b -0.060 0.154 0.686 0.677 0.678
google/gemma-7b -0.059 0.017 0.675 0.544 0.691
GritLM/GritLM-7B -0.125 0.044 0.711 0.702 0.697
HuggingFaceH4/zephyr-7b-beta -0.129 0.097 0.706 0.687 0.699
intfloat/multilingual-e5-large-instruct -0.037 0.096 0.729 0.720 0.724
meta-llama/Llama-2-7b-hf 0.009 -0.032 0.701 0.692 0.686
openai-community/gpt2-large 0.042 -0.038 0.664 0.665 0.669
openai-community/gpt2-xl 0.041 0.071 0.666 0.667 0.669
tiiuae/falcon-7b -0.116 0.066 0.686 0.672 0.694
BAAI/bge-base-en-v1.5 Race 0.094 0.092 0.730 0.709 0.683
BAAI/bge-large-en-v1.5 0.104 0.099 0.727 0.727 0.695
google/gemma-2b -0.041 0.164 0.686 0.665 0.686
google/gemma-7b -0.055 0.133 0.675 0.549 0.678
GritLM/GritLM-7B -0.133 -0.057 0.711 0.714 0.690
HuggingFaceH4/zephyr-7b-beta -0.127 0.062 0.706 0.687 0.697
intfloat/multilingual-e5-large-instruct 0.053 0.120 0.729 0.730 0.730
meta-llama/Llama-2-7b-hf -0.058 0.113 0.701 0.699 0.705
openai-community/gpt2-large -0.019 0.024 0.664 0.670 0.680
openai-community/gpt2-xl 0.149 0.180 0.666 0.665 0.692
tiiuae/falcon-7b -0.192 -0.027 0.686 0.664 0.693
BAAI/bge-base-en-v1.5 Religion 0.054 0.078 0.730 0.716 0.694
BAAI/bge-large-en-v1.5 0.153 0.175 0.727 0.718 0.697
google/gemma-2b 0.118 0.278 0.686 0.679 0.682
google/gemma-7b 0.127 0.194 0.675 0.548 0.685
GritLM/GritLM-7B -0.002 0.077 0.711 0.702 0.703
HuggingFaceH4/zephyr-7b-beta -0.130 0.118 0.706 0.693 0.686
intfloat/multilingual-e5-large-instruct 0.201 0.194 0.729 0.728 0.735
meta-llama/Llama-2-7b-hf -0.103 0.032 0.701 0.679 0.710
openai-community/gpt2-xl 0.247 0.251 0.666 0.671 0.679
tiiuae/falcon-7b 0.126 0.265 0.686 0.671 0.703
Table 7: Downstream testing results on Stanford Sentiment Treebank and STS-B Semantic Similarity Dataset. PCC here refers to the Pearson’s Coefficient and we report the gain in positive PCC from the Biased embeddings to the debiased embeddings. SST is Stanford Sentiment Treebank and STS-B is the Semantic Textual Similarity Benchmark
Appendix E Variation of Bias in the Different LLMs
The presence of biases in has drawn significant attention from researchers and practitioners. These biases can inadvertently emerge during the training process due to the characteristics of the initial training data. In this study, we explore the factors contributing to bias variation among LLMs, focusing on three prominent models: Llama, Mistral, and Gemma. Our analysis reveals that biases, including those related to gender, race, and culture, are often inherited from the training data. For instance, historical texts may perpetuate gender stereotypes or racial prejudices present in their source material. Llama and Mistral, trained on diverse corpora containing web documents, source code, and mathematical text, exhibit varying degrees of bias. Gemma, released by Google, further demonstrates the impact of training data size, with both 2B and 7B variants drawing from an extensive pool of up to 6 trillion tokens.
Appendix F Ablation Experiments
In our study, we conduct ablation experiments to assess the effectiveness of various debiasing techniques in the realm of natural language processing. These techniques encompassed five distinct scenarios: the utilization of debiased embeddings, the application of the original soft debiasing method, the original debiasing method with the Adam optimizer, DeepSoftDebias with the SGD optimizer, and finally, DeepSoftDebias with the Adam optimizer. These experiments were gauged based on MAC as the evaluation metric.
Through rigorous experimentation across three biasing directions, we systematically analyze the performance of each method. Our results reveal a consistent trend of incremental improvements as we transitioned from one method to the next. Notably, DeepSoftDebias, emerged as the standout performer, boasting the highest mean average cosine similarity score across all evaluated scenarios. In addition, our analysis revealed that substituting the transformation matrix with our neural network approach resulted in the most significant enhancement in the efficacy of the debiasing method. This observation underscores the pivotal role played by neural networks in maximizing the effectiveness of the debiasing techniques. Table 8 presents a visualization of the results of our ablation experiments.